OpenStudy (anonymous):

Find the equation of the tangent line y=2/(squareroot) x

OpenStudy (anonymous):

@ApoorvaAnand Help me again? you explain really well!

OpenStudy (anonymous):

y = 2*x^(-1/2) That is the question Now slope of tangent can again be written as, dy/dx = 2*(-1/2)*x^(-3/2) Since we do not know the points, we can't find the eq of tangent. You need to tell us the point at which the eq of tangent should be formed.

OpenStudy (anonymous):

sorry there are points I didnt see them written they are 4 and 2/2

OpenStudy (anonymous):

@ApoorvaAnand

OpenStudy (anonymous):

dy/dx = 2*(-1/2)*(1/8) = -8 now Sub in eq and find your answer (like previous question) :)

OpenStudy (anonymous):

so then do I do 4/(2/2)? @ApoorvaAnand

OpenStudy (anonymous):

No. The eq of tangent is now y-4/x-1 = -8 2/2 = 1

OpenStudy (anonymous):

so then if im solving for y? @ApoorvaAnand

OpenStudy (anonymous):

Yes. Solve for y for y = mx+c form.

OpenStudy (anonymous):

y=-8(4) not sure about the c @ApoorvaAnand

OpenStudy (anonymous):

No. You only substitute the values to y1 and x1. Now y and x. The answer is y-4 = -8x+8 => y = - 8x+12 or the eq of tangent is y + 8x - 12 = 0

OpenStudy (anonymous):

ohh so then y=-8x+12? @ApoorvaAnand

OpenStudy (anonymous):

Yes. :)

OpenStudy (xapproachesinfinity):

is the given point \[\large (4, \frac{2}{2})\] hmmm.....

OpenStudy (xapproachesinfinity):

what is the given point?

OpenStudy (xapproachesinfinity):

the way you wrote it is not clear and apoorva.. used it even though it was not clear

OpenStudy (xapproachesinfinity):

you know post the picture for a better understanding

OpenStudy (xapproachesinfinity):

take a snapshot and post :)

OpenStudy (xapproachesinfinity):

@shaekitchen you should relook at this